### v1

# An engine not a camera: Measuring performative power of online search

Abstract
The power of digital platforms is at the center of major ongoing policy and regulatory efforts. To advance existing debates, we designed and executed an experiment to measure the power of online search providers, building on the recent definition of performative power. Instantiated in our setting, performative power quantifies the ability of a search engine to steer web traffic by rearranging results. To operationalize this definition we developed a browser extension that performs unassuming randomized experiments in the background. These randomized experiments emulate updates to the search algorithm and identify the causal effect of different content arrangements on clicks. We formally relate these causal effects to performative power. Analyzing tens of thousands of clicks, we discuss what our robust quantitative findings say about the power of online search engines. More broadly, we envision our work to serve as a blueprint for how performative power and online experiments can be integrated with future investigations into the economic power of digital platforms.
1 Introduction
At the heart of one of Europe’s most prominent antitrust case is a seemingly mundane question: How much can a search engine redirect traffic through content positioning? In 2017, the European Commission alleged that Google favored its own comparison shopping service by steering clicks away from search results towards Google’s own product comparison service. The technical centerpiece of the case was an ad-hoc data analysis about the position and display biases of Google search results. Google appealed the European Commission’s charges, pointing to, among other arguments, methodological errors. [^2] Case T-612/17, Google and Alphabet v Commission (Google Shopping), ECLI:EU:T:2021:763. The case is still pending, see Case C-48/22 P, Google and Alphabet v Commission (Google Shopping).
The case is emblematic of a broader problem. Although urgently needed, there is currently no accepted technical framework for answering basic questions about the economic power of digital platforms. Lawyers, economists, and policy makers agree that traditional antitrust tools struggle with multi-sided platforms (Stigler Committee, 2019; Crémer et al., 2019). Among other issues, these tools are not able to capture behavioral aspects defining user-platform interactions. Against this backdrop, a recently developed concept, called performative power (Hardt et al., 2022), suggests a way to mitigate limitations of existing antitrust enforcement tools. Performative power measures how much a platform can causally influence platform users through its algorithmic actions. By directly relating power to a causal effect, it sidesteps the complexities underlying conventional market definitions and offers a promising framework to integrate data and experimental methods with digital market investigations. Although the definition of performative power enjoys appealing theoretical properties, a proof of its practical applicability was still missing.
Figure 1: The ability to influence web traffic through content arrangement. Blue bars show average click probability observed for generic search results in position 1 to 6 on Google search under different counterfactual arrangements; default arrangement (left), swapping results 1 and 2 (middle), swapping results 1 and 3 (right). We see a significant effect of these arrangement changes on clicks. We will inspect this more formally in Section 5 where we also explore arrangement changes beyond reranking.
1.1 Our work
In this work, we present a first proof of concept showing how to estimate performative power of a search engine in practice. The instantiation of performative power we are interested in concerns the ability of a search engine to impact web traffic through decisions of how to arrange content.
Our core contribution is to design and implement an online experiment to provide quantitative insights into the causal effect of content arrangement on clicks for the two most widely used search engines, Google Search and Bing. Our experiment is based on a browser extension, called Powermeter, that emulates small updates to the platform’s algorithm by modifying how search results are displayed to users. We implement different counterfactual arrangements to inspect the effect of re-ranking and favored positioning (e.g., Ads or Shopping boxes) on clicks, both in isolation and jointly; the latter is relevant for the Google Shopping investigation. The arrangement to which a user is exposed is chosen at random every time they perform a search. We discuss several technical steps we implemented to take care of the internal validity of our experimental design.
Using Powermeter we collected data of about 57,000 search queries from more than 80 different subjects, over the period of 5 months. The queries for any given user follow the distribution of queries under their every-day use of online search. Our data allows us to measure the causal effects of arrangement under natural interactions of users with the platform. Figure 1 provides a first glimpse into the observed effects. In summary, we find that consistently down-ranking the first element by one position causes a reduction in clicks of $42\%$ for the respective element on Google search on average. Down-ranking the same element by two positions yields a reduction of more than $50\%$. For Bing we find an even larger effect of ranking, although with less tight confidence intervals due to the small number of Bing queries performed by our participants. Inspecting different subsets of queries we find that the effect of position is larger for queries with a high number of candidate search results. When combining down-ranking with the addition of Ads or Shopping boxes, the effect of arrangement is even more pronounced, showing a distortion in click for the first result of $66\%$ averaged across queries where such elements are naturally present on Google search. To the best of our knowledge, we are the first to publish experimental insights into the display effects on Google search and Bing.
Finally, we outline how to relate our measurements to performative power relevant in the context of digital market investigations. We take a recent antitrust investigation against Alphabet Inc. ran by the European Commission as a case study. Taken together, we hope our empirical and theoretical results can serve as a first blueprint for what future antitrust investigations of digital platforms’ market power might look like.
2 Background and related work
The market power of digital platforms is the subject of a robust debate in policy, legal, and technical circles. See, for example, Newman (2014); Crémer et al. (2019); Stigler Committee (2019); Furman (2019); Cabral et al. (2021). Conventional antitrust enforcement tools have been put into question (Syverson, 2019) and new concepts of market power have been proposed to deal with the complexities of digital markets (OECD, 2022). These account for the multi-sided nature of the markets as well as the role of behavioral weaknesses of consumers—albeit with limited success. We refer to a comprehensive literature survey about behavioral aspects in online market by the UK Competition and Market Authority (CMA, 2017).
Performative power.
Inspired by recent developments in the computer science literature, Hardt et al. (2022) proposed performative power as a measure of influence that predictive systems can have over participants. The term performativity more broadly refers to a causal effect of predictions on the target of prediction. The study of performativity in economics dates back to Morgenstern (1928); Grunberg and Modigliani (1954); Simon (1954). Performative power builds on the more recent development of performative prediction (Perdomo et al., 2020). Performative prediction introduces the concept of performativity to machine learning and exposes its role in data-driven predictions. We refer the reader to Hardt and Mendler-Dünner (2023) for a recent survey on the topic. A robust insight from performative prediction is that there are two ways to succeed at prediction. One is to learn patterns in the data-generating distribution. The other is to steer the data-generating distribution to a desirable state. Performative power (Hardt et al., 2022) builds on these insights and starts from a simple observation. The degree to which predictions are performative depends on the power—in terms of reach and scale—of the platform making the predictions. Thus, by measuring the degree to which predictions are performative, we can estimate the power of a platform. Formally, performative power relates the ability of a platform to steer the population of participants through predictions, to the causal effect of algorithmic actions. We provide the definition below.
Definition 1 (Performative power (Hardt et al., 2022)).
Given the algorithmic action $a_{0}$ and a set of alternative conducts $\mathcal{A}$, a population $\mathcal{Q}$ and an outcome variable $z$. Performative power is defined as
$\mathrm{PP}:=\sup_{a\in\mathcal{A}}\;\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\operatorname*{\mathbb{E}}\|z_{a_{0}}(q)-z_{a}(q)\|_{1}\,,$ (1)
where $z_{a_{0}}(q)$ denotes the outcome for unit $q\in\mathcal{Q}$ under $a_{0}$ and $z_{a}(q)$ denotes the counterfactual outcome, would the platform implement $a\in\mathcal{A}$ instead. Expectations are taken with respect to the randomness in the potential outcome.
In words, performative power measures the ability of a firm to cause a change in the outcome on average across a population, by means of changing algorithmic actions. As such, performative power offers a family of definitions that can be instantiated flexibly in a given context. The distance notion could also be generalized, which we refrain from for the sake of presentation. For this work we discuss a concrete instantiation of performative power motivated by a recent antitrust investigation against Alphabet Inc. ran by the European Commission, [^3] European Commission, AT.39740, Google Search (Shopping), 27.06.2017. referred to as the Google Shopping case.
Google shopping case.
The subject of the investigation is the alleged conduct on Google search of introducing a) visually more appealing boxes for product comparisons on shopping queries in combination with b) an update to the ranking algorithm that leads to competing product comparison services appearing further down in the generic search results. The combination of these two conducts is made responsible for distorting outgoing web traffic from general search, and harm competition in the related specialised search markets (e.g. shopping). We are interested in providing empirical evidence into this effect.
We build on a result of Hardt et al. (2022) that formally relate performative power to the causal effect of position. For our investigation, we generalize the effect of position to encompassing algorithmic updates beyond reranking. These updates can be seen as changes to what is commonly referred to as choice architecture (Thaler and Sunstein, 2008) controlled by the platform’s algorithm. We focus on online search. This is worth emphasizing, as the causal effect of algorithmic changes is a result of users’ on-platform behavior and influenced by behavioral weaknesses, biases, and preferences. Thus, its strength strongly depends on the context in which users interact with the platform. To account for this we perform our experiments and measurements under natural interactions of users with the platform.
Causal effect of position.
Evidence for position bias was first reported in eye-tracking studies by Joachims et al. (2005); Guan and Cutrell (2007). Since then the importance of position bias in recommender systems more broadly, has been studied in multiple contexts. It play an important role when interpreting user feedback in recommender systems (Joachims et al., 2005), designing auctions (Aggarwal et al., 2006; Athey and Ellison, 2011), learning-to-rank (Agarwal et al., 2019; Qin et al., 2020; Chen et al., 2022; Zhang et al., 2023), debiasing click data (Craswell et al., 2008; Chuklin and and, 2015; Xv et al., 2022; Chen et al., 2023), and understanding incentives on different sides of the market (Ben-Porat and Tennenholtz, 2018; Hron et al., 2023; Hardt et al., 2023), to name a few. Most directly related to our work are efforts that aim to measure position bias: Ursu (2018) rely on public data collected under randomized result ordering to investigate the role of positioning on Expedia. Narayanan and Kalyanam (2015) investigated position bias in search advertising using a regression discontinuity design. Also focusing on online advertising, Agarwal et al. (2011) investigate position bias by experimentally randomizing bids to influence the ranking. With our work we demonstrate how a browser extension can serve as a useful tool for directly measuring the causal effect of position without having access to the platform’s algorithm. This allows us to provide experimental insights into multiple algorithmic changes. Browser extensions have previously been used for automatically collecting data to audit systems. Robertson et al. (2018) audit Google search for polarization on politically-related searches. Gleason et al. (2023) collect data via an extension to directly investigate the effect of search result components on clicks. Also the ongoing National Internet Observatory (NSF, 2021) relies on a browser extension to collect web traffic data. While most prior works focus on observational data, we use the extension to conduct online experiments.
3 Performativity in online search
We set up the causal question under investigation. Therefore, we model an online search platform as a distribution over events. An event is a triplet of a user query $Q$, content arrangement $A$ and click outcome $C$. A user query corresponds to a person visiting the search page and entering a search query in the search bar. The query is processed by the platform and results in an arrangement of content on the website. The mapping is typically defined by a proprietary pipeline involving a ranking algorithm that determines the order in which search results are ranked and displayed, including the positioning of components such as Ads or featured elements. Then, mediated by the arrangement, the user query leads to a click outcome $C$. The categorical random variable $C$ indexes the element clicked over by the user. It is a function of the user query and the arrangement.
3.1 The causal effect of arrangement
Assume the platforms were to change their algorithm that determines the content arrangement. We seek to answer the following causal questions: How much does a change to the arrangement impact clicks of a content element on the search page?
If clicks were solely determined by stable preferences, then we would see no effect. Performativity is the reason why we see an effect. Behavioral biases, weaknesses, the limited ability to process large amounts of data, and trust in the platform can be a source for performativity. The more performative the arrangement is, the stronger the effect. We use $a_{0}$ to refer to the reference arrangement of results on Google search. For a given user query $q$ we define the potential outcome of a click event under the arrangement $a$ as $C_{q}(a)$. The variable $C$ takes on categorical values, indexing the elements on the page. We mainly focus on clicks to general search results. Let $\{c_{1},...c_{K}\}$ denote the top $K$ general search results indexed in the order they appear under $a_{0}$.
Definition 2 (Performativity gap).
Given a counterfactual arrangement $a^{\prime}$, we define the performativity gap at position $i$ with respect to a population of queries $\mathcal{Q}$ as
$\mathrm{\delta}^{i}(a^{\prime})=\operatorname*{\mathbb{E}}\big{[}1\{C_{q}(a^{\prime})=c_{i}\}\big{]}-\operatorname*{\mathbb{E}}\big{[}1\{C_{q}(a_{0})=c_{i}\}\big{]}\,,$ (2)
where expectations are taken over $q\in\mathcal{Q}$ and the randomness in the potential outcome.
In words, the performativity gap quantifies how much the expected click through rate of search item $c_{i}$ on average across queries $\mathcal{Q}$ changed, had the platform deployed arrangement $a^{\prime}$ instead of arrangement $a_{0}$. We are particularly interested in counterfactuals $a^{\prime}$ that constitute small changes to the original arrangement $a_{0}$, such as the slight reordering of search results, or the addition of visually appealing elements on the website. The performativity gap offers a generalization of the causal effect of position defined in Hardt et al. (2022), to cover algorithmic updates beyond reranking. We will discuss concrete instantiations in Section 4.2.
More broadly, the effect of small arrangement changes provides a lower bound on the ability of a platform to impact clicks with potentially more significant algorithmic changes. The following result generalizing Theorem 8 in Hardt et al. (2022) is a direct consequence:
Theorem 1 (Generalization of Hardt et al. (2022); Theorem 8).
Let $\mathrm{PP}$ be the performative power of a search platform defined with respect to a set of arrangements $\mathcal{A}$, a population of search queries $\mathcal{Q}$ performed on the platform, and the outcome variable $z_{a}(q)=\mathrm{1}\{C_{q}(a)=c_{1}\}$. Then, performative power is lower bounded by the performativity gap as $\mathrm{PP}\geq\sup_{a\in\mathcal{A}}\delta^{1}(a)$.
Algorithmic distortion.
Often it can be useful to express the performativity gap relative to the base click rate. Thus, we define the algorithmic distortion factor as the smallest factor $\beta>0$ such that
$\mathrm{\delta}^{i}(a^{\prime})\leq\beta\operatorname*{\mathbb{E}}\big{[}1\{C_{q}(a_{0})=c_{i}\}\big{]}.$ (3)
This quantity serves as a way to denote the fraction of clicks taken away from content item $c_{i}$ as a result of the update $a_{0}\rightarrow a^{\prime}$. Also, it helps to express power relative to a base click through rate as it offers a more interpretable quantity for investigators.
3.2 Estimation using an RCT
To estimate the performativity gap, we rely on a randomized controlled trial (RCT), the gold standard methodology to estimate causal effects (Fisher, 1935; Imbens and Rubin, 2015). Given the fundamental challenge of causal inference that we can not observe a search query simultaneously exposed to different arrangements, the idea of an RCT is to randomly select, for each query $q\in\mathcal{Q}$, the arrangement they are exposed to. We write $Q_{a}\subset\mathcal{Q}$ for the subset of queries that are exposed to treatment $A=a$. We also refer these subsets as treatment groups. Observing the click events for each queries allows us to obtain an estimate of the performativity gap as
${\mathrm{\bar{\delta}}}^{i}(a^{\prime})=\mathrm{{CTR}}^{i}(a^{\prime})-\mathrm{{CTR}}^{i}(a_{0})\quad\text{with}\quad\mathrm{CTR}^{i}(a)=\frac{1}{|Q_{a}|}\sum_{q\in Q_{a}}\mathrm{1}\{C_{q}(a)=c_{i}\}.$
For this ${\mathrm{\bar{\delta}}}^{i}(a^{\prime})$ to provide an unbiased estimate of $\delta^{i}(a^{\prime})$, we rely on the following application of the stable unit treatment value assumption (SUTVA) (Rubin, 1980) to be satisfied:
Assumption 1 (Independence across queries).
[…]
In the following section we discuss steps we take in our experimental design towards justifying this Assumption 1. Beyond requiring relatively weak assumptions, the advantage of using an experimental approach to measure the performativity gap is that, while the mechanism mapping user queries to clicks can be arbitrarily complex, this complexity does not affect the experiment. Aspects such as users’ preference for clicking links on the left side of the screen (Dimitrov et al., 2016), users’ trust in the platform ranking (Pan et al., 2007), the effect of visually appealing elements (Fubel et al., 2023), and other behavioral biases will naturally enter our measurement. We directly measure their effect on the relevant quantity of interest, namely, the resulting clicks.
4 Powermeter: Experiment in the wild
We designed an online experiment to measure the performativity gap on two popular online search platforms operated by Google and Bing. The experiment is built around a Chrome browser extension that modifies the arrangement of search result pages and records user click statistics in a privacy-preserving fashion. Thereby, the extension allows us to estimate the performativity gap over an organic set of clicks and arrangements without having control over the platforms’ underlying algorithm.
4.1 Browser extension
Powermeter is a Chrome browser extension designed to implement different counterfactual arrangements on Google search and Bing search. Extensions can add functionalities to the web-browser and change how a website is displayed to the user. We emphasize that Powermeter takes advantage of these capabilities, to hide or reorder, but never modify or add any content on the search page.
Technical details.
Once activated, the extension triggers the experiment whenever the user enters a search query on either Google search or Bing search. This can be identified by monitoring the url string of the current tab. Before search results are loaded the extension immediately hides the content of the website, inspects the html document, randomly assigns the user to one of the experimental groups and then implements the respective changes to the website before making the page visible. The implementation of the counterfactuals is done by identifying the relevant items to hide or swap by their html class names or ids. We also add custom tags and event listeners to the identified elements that we can fall back on at a later stage. The entire setup of the experiment usually takes around 40 milliseconds. Hiding the html body of the website with the first possible Chrome event is crucial to avoid glitches in case of bad internet connection and make sure the control arrangement is not revealed to the participant. To ensure internal validity of our experiment, we also have to ensure a participant is not reassigned to a new experimental group when reloading a page, navigating between tabs or repeatedly entering the same search query. This is done by storing a hash of user ID and search query together with the assigned experimental group in the browser cache.
Backend and data collection.
Every participant is assigned a unique random number that serves as anonymous user ID upon installation of the extension. This user ID is stored in the local browser cache and persists throughout the experiment. Every time a click event on an element on the search page is registered, the click data is aggregated into a json object and sent to a database server hosted locally at out institution via a post request using the encrypted https protocol. This concerns information about the index of the clicked search result, the click element type, the page index, and the experimental group. In addition, statistics about the website such as the number of search results, the presence of ads and boxes, the number of candidate results, and the position of specialized search results are extracted from the website are recorded. The database server is built using the Microsoft .Net core framework and deployed within a docker container. The database access is rate limited and the Get endpoint of the database is key protected. We use a SQlite database that is mapped to persistent memory.
Privacy considerations.
The information that is stored with every click does not contain any personally identifiable information. While we record statistics and the position of the clicked element on the search page, we never store search queries or any information about the websites visited by the user. This is an intentional choice to preserve user privacy, and demonstrate that valuable insights can be gathered without privacy invasive data collection. The experiment went through an internal approval procedure and the privacy policy can be found on our website. [^4] The project website can be found at: https://powermeter.is.tue.mpg.de/ We also provide a copy in the appendix.
4.2 Experimental groups
We implement six different counterfactual arrangements, summarized in Table 1. We refer to Figure 2 for the terminology used to refer to individual elements on the general search page. The first three arrangements $a_{1}$$a_{3}$ concern the reordering of organic search results, leaving the other elements on the website untouched. The arrangements $a_{4}$ and $a_{6}$ perform modifications not directly concerning organic search results; Arrangement $a_{6}$ hides a specific element, called the Shopping box, appearing either in the right side panel or on top of the search results page, Arrangement $a_{4}$ hides the box together with all the Ads. Finally, Arrangement $a_{5}$ combines the latter change with a change in search result order. For Bing we only implemented the counterfactual $a_{1}$ to ensure statistical power despite data scarcity. A practical reason not to implement larger modifications is also users’ sensitivity to the resulting deterioration of quality. The Bing experiment of the European Commissions’s investigation had to be discontinued after one week for that exact reason. [^5] See Case T-612/17, supra 2, para 399 Furthermore, our minimal modifications are unlikely to have an immediate effect on user’s browsing behavior, or even be perceived by the user, which is important towards ensuring Assumption 1.
Figure 2: Illustration of different elements on the Google search website.
Table 1: Counterfactual content arrangements implemented by Powermeter as part of the RCT.
Arrangement Description
$a_{0}$ control Search results are displayed without any modification.
$a_{1}$ swap 1-2 The position of the first and the second generic search result are swapped.
$a_{2}$ swap 1-3 The position of the first and the third generic search result are swapped.
$a_{3}$ swap 2-3 The position of the second and the third generic search result are swapped.
$a_{4}$ hide Ads/Box Top Ads and Shopping boxes are hidden.
$a_{5}$ hide + swap Combines the latter modification ($a_{4}$) with swap 1-2 ($a_{1}$).
$a_{6}$ hide Box The shopping boxes are hidden.
4.3 Onboarding
Participants were provided the link to the project website as an entry point. The website contains information about the experiment, the purpose of the study, an onboarding video, as well as the privacy policy of the extension. The extension itself is distributed through the official Chrome webstore and there is a button directly navigating the user to the store. We did not list the extension publicly to ensure participants are informed about the purpose of the study, and protect the integrity of our data. The installation follows the standard procedure of adding a browser extension to Chrome. The user has to give consent to access Google and Bing websites, as well as to use the storage API. The extension remains active until participants remove it from their web-browser, or until the experiment is stopped. The study participants are trusted individuals of different age groups and backgrounds, recruited by reaching out personally or via email. For each of the participants we observe all the search queries they perform while the extension is activated. We provide demographic statistics over our pool of participants in Figure 9 in the appendix.
Data preprocessing.
For each participant we ignored the clicks collected during the first four days after onboarding in order to avoid confounding by changes in behavior due to participation awareness (Keusch et al., 2022). We also removed clicks where the search elements could not be identified reliably for implementing the RCT to avoid selection bias towards the control group.
5 Empirical results
Using our Powermeter browser extension we collected click data from $85$ participants over the course of 4 months, from September 2023 until January 2024. This resulted in $56,971$ click events, and a total of $45,625$ clicks after preprocessing. Out of the clicks $98.9\%$ were registered on Google, and $1.1\%$ on Bing. Figure 8 in the appendix visualizes some aggregate statistics over the clicks collected. We will consider several subsets of these events for which we measure the performativity gap and algorithmic distortion. In the following we discuss the main insights from the collected data. For all plots, we provide bootstrap confidence bounds over 200 resamples.
5.1 Reordering search results
We first inspect the three counterfactual arrangements $a_{1}$, $a_{2}$, $a_{3}$. Recall that $c_{i}$ indexes search result $i$ in the order in which it appears under $a_{0}$. In Figure 3 we visualize the event probabilities $C=c_{i}$ for the control group (blue bars) and compare it to the respective probabilities under the three counterfactuals (orange bars). The three figures on the left show the results across Google search queries. Here the counterfactuals correspond to swapping the position of the first two results (left), the first and the third (middle) and the second and the third (right). The right figure shows the results evaluated on Bing search queries when the first two results are swapped. The lower figure visualizes the corresponding performativity gap $\delta^{i}(a)$ for $i=1,2,...,6$.
We observe a consistently large effect of arrangement on clicks. Being down ranked by one position on Google search decreases average click through rate of $c_{1}$ from $43\%$ to $24\%$, resulting in $\delta^{1}(a_{1})=-0.19$ and an algorithmic distortion of $\beta=0.44$. Being down-ranked by 2 positions results in $\delta^{1}(a_{2})=-0.27$ and an average loss of more than $50\%$ of traffic. Note that a similar effect size has been reported in the case decision for the UK market for a two position shift (European Commission, 2017, 460), indicating that the estimate is robust across different populations. Finally, by down-ranking the second content element by one position we still observe a significant traffic reduction, corresponding to an algorithmic distortion of $\beta=0.39$.
An interesting observation is also that for every counterfactual arrangement, the element shown first ends up getting most clicks on average, despite implementing a different ranking. Implying that all the rankings are performatively stable (Perdomo et al., 2020) with respect to the non-personalized reranking strategies considered. However, there are several indications of Google’s ranking $a_{0}$ reflecting relevance of search results better then the other arrangements. Namely, $c_{1}$ gets more clicks when displayed first, compared to other results in the same position, e.g., the first result under $a_{2}$ or $a_{3}$. A second indicator is that $c_{2}$ benefits from arrangement $a_{3}$; under the hypothesis that users consider search results in order this indicates that content item $c_{3}$ absorbs less clicks than $c_{1}$.
Figure 3: Click through rate and performativity gap under the counterfactual arrangements $a_{1}$, $a_{2}$, $a_{3}$ for Google and the counterfactual arrangement $a_{1}$ for Bing, compared to the control arrangement $a_{0}$.
For Bing the position effect seems to be even more pronounced, although confidence intervals are significantly larger in this case. We conjecture that this could be due to the average number of specialized search results and Ads present on the search page being larger on Bing, as this results in a larger spacial separation of search results. Statistics about elements present on the the search pages shown to our study participants are reported in Figure 8 in the Appendix.
5.2 Indirect effect of visually appealing elements
Next, consider the counterfactuals $a_{4}$ and $a_{6}$ that leave generic search results untouched and hide certain elements on the website. We first inspect the number of clicks these elements absorb if added to the page. We focus on Google search. In Figure 4 we compare the fraction of clicks going to generic search results, Ads, and Boxes for $a_{0}$, $a_{4}$, $a_{6}$. We plot the statistics across the aggregate queries (left panel), the subset of queries where Ads are present on the page under $a_{0}$ (middle panel) and the subset of queries where the box is present on the page under $a_{0}$ (right panel). We find the addition of boxes absorbs $22.4\%$ of the clicks on average across queries where it is present under $a_{0}$ and these clicks are mostly taken away form the generic search results. Similarly, Ads absorb close to $30\%$ of the clicks on average for queries where they are present. However, considering the overall number of clicks the effect is smaller since a large fraction of queries contains neither Ads nor Boxes.
Figure 4: Effect of arrangement on the click distribution across different element types (generic search results, Ads, boxes), visualized for three different subsets of Google queries.
Conduct Distortion
of $c_{1}$
add Box $\beta=0.23$
add Ads/Box $\beta=0.44$
swap+add Box $\beta=0.53$
swap+add Ads/Box $\beta=0.66$
Figure 5: Effect of hiding box and swapping elements on the click probability of generic search results. Statistics are evaluated for the subset of queries for which box is naturally present on Google. The hashed bar shows the click probability under $a_{5}$ when top content is hidden and the first two elements are swapped. The right table reports the empirical measure of algorithmic distortion for different conducts, extracted form the results in the left figure.
Combined conduct.
We now consider the combined conduct of adding the box and down-ranking an element. We constrain our focus on queries where the Shopping box is present under $a_{0}$, either in the center column or in the right sidebar. These are $3.2\%$ of all the events.
We show the corresponding click probabilities for the three first search results in Figure 5. In both figures the blue bars correspond to the control group $a_{0}$, and the red bars correspond to $a_{1}$. For these groups boxes are present on the page. In the left figure we investigate the effect of hiding boxes only and the orange bars correspond to arrangement $a_{4}$. In the right figure we investigate the effect of also hiding Ads, here the orange bars correspond to $a_{6}$. We find that when adding Boxes, all three content items loose a significant fraction of clicks, whereas Ads mostly take away from $c_{1}$. In the right figure we additionally show $a_{5}$ using the hatched bars (i.e., down-ranking the first element by one position if box is hidden). For $c_{1}$ the combined effect of adding Ads and Boxes on the click through rate is almost as large as the effect of down ranking the same item by one position. What we can see consistently is that combining the conduct of adding visually appealing elements on top of the page, and down-ranking a content item, has a combined effect that is larger than the effect of the two individual modifications alone. For element $c_{1}$ the measured distortion is reported in the table. The combined effect leads to a reduction of $25\%$ in clicks and an algorithmic distortion of $66\%$ when considering the effect of Ads/Boxes, and $53\%$ when only considering Boxes (comparing orange and red bar).
5.3 Factors that impact performativity
We perform additional investigations into what factors have a reinforcing effect on the performativity of the ranking. To this end, we inspect different subsets of Google queries and measure the performativity gap as well as algorithmic distortion for $c_{1}$ under the counterfactual $a_{1}$. First, we split the data across two different axes depending on whether Ads or boxes are present, and whether Specialized Search results (SSR) are present in between the first two results. The respective comparisons are visualized in the left and middle panel of Figure 6. We observe that the performativity gap in the presence of Ads and boxes is smaller, and about the same whether special search results are present in between search results. However, if we plot distortion we get a different picture, since the base click probability of content item $c_{1}$ across different splits can be very different for the three cases. We find that while Ads and Boxes have little effect on distortion, the presence of a specialized search result in between the swapped results tends to increase the effect of down-ranking on $c_{1}$.
Figure 6: Performativity gap and algorithmic distortion for content item $c_{1}$ under the counterfactual arrangement $a_{1}$ measure across different subsets of Google search queries.
For the second investigation we group the queries by the number of candidate search results available on Google. This number was extracted form the website where it appeared as a string on top of the page in the form ‘About 323’000’000 results (0.65 second)’. The right panel of Figure 6 shows algorithmic distortion for each percentile of the data. We see a clear trend that the performativity gap increases with the number of candidate results. We suspect this to be connected to the smaller relevance gap across results for queries with more potential results, leading to a higher influence of the arrangement on clicks. However, note that findings in this subsection are no causal claims.
6 Discussion
In this work we present a flexible experimental design, based on a browser extension, to investigate the effect of search algorithms on user clicks. The browser extension performs interventions at the level of display to emulate algorithmic updates, without access to the platform algorithm. We implement different counterfactuals, and measure their effect on clicks. In turn, we formally relate our quantitative findings to an instantiation of performative power measuring the platform’s ability to steer user clicks going out from its website.
Next, we describe how our findings could fit into a broader anti-trust investigation. Take the Google Shopping case as an example. It is concerned with the ability of Google search to distort incoming traffic to a particular business’ website. Suppose, as a thought experiment, $80\%$ of the search referrals to the site come from Google Search. [^6] As part of the Google Shopping investigation (Section 7.2.4.1 in the decision, Table 24) it is reported that up to $82\%$ of incoming traffic to CS services in the European Economic Area (EEA) was mediated by Google search in 2012. Further, assume that $70\%$ of the referrals from Google happen while the service is ranked among the top two generic search results. Our estimates suggest that distortion of traffic at the first position can be as large as $66\%$ for small arrangement changes. Assuming for the second position the effect is $20\%$ smaller, giving a conservative average effect size of $\beta=0.8\cdot 0.66$. Multiplying the effect size by the fraction of incoming clicks it concerns, we get $0.8\cdot 0.7\cdot 0.9\cdot 0.66\approx 33\%$. This is the fraction of traffic to the site Google can influence.
Turning this number into a conservative lower bound on performative power it offers an interpretable measure of power for an investigator to judge whether the use of this algorithmic lever should be a concern for competition and consumers or not. We can use the same logic to compare search engines, and assess the effectiveness of mandated remedies. One avenue for future work to strengthen such an argument is to rigorously argue about the external validity of empirical findings like ours.
More broadly, we hope our work can serve as a blue print for how performative power can be used to guide digital market investigations, and how tools from computer science, causality, and performative prediction can inform ongoing legal debates related to the power of digital platforms. Our work is situated within a growing scholarship (c.f., Metaxa et al., 2021; Bandy, 2021; Dash et al., 2024) that takes advantage of the accessibility of digital markets for monitoring and regulating digital platforms. In addition to observational methods, we demonstrate how experiments can offer an additional powerful lever.
From the perspective of computer science our work offers measurements and provides evidence of performativity in the context of online search, contributing empirical support to the study of performativity on online platforms. As its name suggests, a search engine acts as an engine steering consumption through its ranking algorithm, rather than a camera merely picturing candidate results—we borrow this analogy from MacKenzie (2008).
Acknowledgements
We are particularly grateful to everyone who installed our extension and participated in the study; without you such a project would not have been feasible. Further, the authors would like to thank Jonathan Williams for the design of the project website, Alejandro Posada for producing the onboarding video and helping with the logo, Telintor Ntounis for assistance in setting up the server infrastructure, and Ana-Andreea Stoica, André Cruz and Jiduan Wu for feedback on the manuscript.
Appendix A Causal model
To support the discussion of our causal claim, consider the simplified diagram of how a user navigates to a website illustrated in Figure 7. The random variable $U$ denotes a user request and $T$ indicates which website is being visited in response. The user either navigates to the website via Google search (gray box), or they navigate to the website on some alternative way. This can be by entering the url directly, or using a different search service. Naturally, the arrangement $A$ only impacts the outcome $T$ if the user uses Google search, otherwise $A$ does not have any influence on the outcome. For a search query, the user query leads to an arrangement of content shown on the website, and the arrangement mediates the click. The setup in Section 3 focuses on performative power within the grey box, and Section 6 extends the claims to the full graph.
[…]
Figure 7: Simplified causal graph illustrating the use of an online service. A web request leads to the visit of a website, potentially mediated by Google search.
Appendix B Proof of Theorem 1
We instantiate performative power with respect to a set of action $\mathcal{A}$, a population of search queries $\mathcal{Q}$, and the outcome variable $z_{a}(q)=\mathrm{1}[C_{q}(a)=c_{i}]$. Since the outcome is a scalar the $L_{1}$ norm reduces to an absolute value and we get
$\mathrm{PP}=\sup_{a\in\mathcal{A}}\;\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\operatorname*{\mathbb{E}}|z_{a_{0}}(q)-z_{a}(q)|$ (4)
Using the definition of the performativity gap and the definition of the $L_{1}$ norm we have that
$\mathrm{PP}$ $=\sup_{a\in\mathcal{A}}\;\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\operatorname*{\mathbb{E}}|\mathrm{1}[C_{q}(a_{0})=c_{i}]-\mathrm{1}[C_{q}(a)=c_{i}]|$
$\geq\sup_{a\in\mathcal{A}}\;\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\left|\mathrm{Pr}[C_{q}(a_{0})=c_{i}]-\mathrm{Pr}[C_{q}(a)=c_{i}]\right|$
$\geq\sup_{a\in\mathcal{A}}\;\Big{|}\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathrm{Pr}[C_{q}(a_{0})=c_{i}]-\frac{1}{|\mathcal{Q}|}\sum_{q\in\mathcal{Q}}\mathrm{Pr}[C_{q}(a)=c_{i}]\Big{|}$
$=\sup_{a\in\mathcal{A}}\delta^{i}(a)$
where the performativity gap is defined with respect to the set of queries $\mathcal{Q}$.
Appendix C Additional details on the study
C.1 Aggregate click statistics
In Figure 8 we show aggregate statistics over the clicks collected. In Figure 9 we provide aggregate statistics over the user base, collected through the onboarding form.
Figure 8: Aggregate statistics over clicks and search result pages collected during our experiment. The blue bars show the statistics for Google and the orange bars show the statistics for Bing. Numbers are aggregated based on original search pages, before any modifications are performed.
Figure 9: Aggregate user statistics collected from the 85 participants with the onboarding form. Age distribution (left) and language in which they consume online search (right). This is the data we have about the demographics of our participants.
C.2 On the external validity of the Powermeter design
In our work we have demonstrated the large effect of arrangements on clicks for the participants of our study. We also discussed the robustness of our quantitative findings to variations in the sample of queries using bootstrapped confidence intervals.
Now discuss how the quantitative findings from the Powermeter experiment can be extrapolated to claims about a broader population. The key step is to argue that the observed sample of search queries is sufficiently representative for the population of interest. To this end, we note that the Powermeter design has the intrinsic property that the observed queries of an individual users follows their natural interaction with the platform. In turn, the representativeness of the user base has to be considered on a case-by-case basis, depending on the population under investigation. In technical terms, our sample of queries corresponds to a single-stage cluster sample of $\mathcal{Q}$. This is a commonly encountered sampling methodology in human subject studies.
Unbiasedness.
Our cluster sample estimates are unbiased, if the sample of users is also unbiased. Propensity score reweighting Stuart et al. (2011) can help to adjust estimates. To support such adjustments, Powermeter collects minimal demographic statistics about the users (see Figure 9). Further, this data is linked with collected clicks so the data can be reweighted appropriately.
Statistical error bounds.
For assessing the statistical error bounds of empirical estimates obtained with cluster sampling, it has to be taken into account to what extent the sample $S$ represent the variations in the population $\mathcal{Q}$. See (e.g., Rutterford et al., 2015; Abadie et al., 2022) for a reference. At the high level, assume there are $k$ types of users that respond differently to an arrangement. If all these user types are proportionally represented in the sample $S$, then the statistical uncertainty of our measurements scales with the number of queries as
$\delta^{1}(a)\geq\delta_{S}^{1}(a)-\mathcal{O}(1/\sqrt{|Q_{S}|}).$ (5)
In turn, if there is potential variation in the sampling of these types this also need to be taken into account. As long as variations across search queries dominate variations across users, the number of queries dominates the errors of the Powermeter estimates, but a correction term in (5) is necessary. Formally establishing these trade-offs, is an interesting step for future work.
Related to our findings, we conjecture that in the context of online search intra-cluster variation is more pronounced than inter-cluster variation. Meaning that the effect of position varies more across different type of search queries (whether it is an exploratory search for a suitable restaurant in a remote city, or a search for a specific research paper) than it does across users. This assessment might change is the consideration is restricted to specific types of queries.
C.3 Project website
The project website (https://powermeter.is.tue.mpg.de/) provides the entry point to the experiment. On the website users were provided with details about the experiment, instructions for how to participate, how to install the extension and information about how their data is used. We provide a list of frequently asked questions, as well as a instruction video. See Figure 10 for a screenshot. When clicking ‘participate’ the user’ is navigated to the extension in the Chrome store. We did not list the extension publicly to make sure the users have visited the project website before participating.
C.4 Onboarding form
Upon installation of the extension the user is automatically navigated to the onboarding form, as illustrated in Figure 11. Filling the form is voluntary and answers are binned to only provide coarse grained information and no personally identifiable information. The information serves to gather some statistics about the demographics of the participants, but also to potentially debug different languages and website versions. Typically coarse-grained information about the statistics is sufficient to argue about qualitative properties of the population, e.g., for the sake of propensity adjustment.
C.5 Code
The extension is implemented in Java Script using the Google Manifest V3 framework. The extension is available through the Google Chrome store. The current version of the extension is v1.3. The experiment in this version is deactivated through a flag in the code. It is no longer maintained to keep up with recent updates on Google search. However, it is still available and the source code can be inspected using the Chrome DevTools.
Figure 10: Project website. Screencapture form May 29, 2024.
Figure 11: Onboarding form. Screencapture form May 29, 2024.
C.6 Extension policy
Below we provide the content of our extension policy. It is available through the Chrome store and published on our website:
[…] are committed to protect your data and to respect your privacy. This privacy statement explains the reason for the collection and processing of your data in the context of the Powermeter Research Experiment. It explains what data is collected, the way in which it is collected, handled, and protected, how that information is used.
Why and how do we process your data?
[…] collects and further processes your data to conduct the Powermeter Research Experiment (before, during and after) and to process your registration in that experiment. Powermeter is a Chrome web-browser extension (i.e. a software application adding functionality to a web browser) designed exclusively for research purposes and distributed through the Google Chrome Store. By installing the extension, users agree to participate in a research experiment. Powermeter performs random modifications to the display or positioning of search results in popular search engines to investigate the causal effect of display and positioning on user click behavior. Powermeter does not modify individual search results or links. It only rearranges ranking, positioning, and display of search results. It does not change, modify, distribute, or otherwise alter any part of the search engines’ software it operates on.
Powermeter does not collect personal data. It indexes users within the population through installation time-stamps and then counts the number of clicks occurring on search engines associated with different arrangements of content as inputted by each user. The indexing of users is necessary to understand if clicks are coming from different or from the same user and for the evaluation of statistics across different demographics. Specifically, Powermeter collects the following non-personal data: (i) the type of search engine used (Google or Bing); (iii) the version of the extension installed; (iii) the index of the search page where the click occurs; (iv) the user index; (v) the number and relative positioning of ads and specialised search results present on the page the click occurs; (vi) the relative positioning of the clicked item vis-à-vis related items object of interest; (vii) the timestamp of when the click occurred.
Beyond that, Powermeter never stores users’ search query, users’ identifiers (such as names, emails, IP addresses, etc.), nor records any information regarding the content of the links clicked over by users or the url of the landing page or any other data related to users’ browsing activity. The use of information received from Google APIs will adhere to the Chrome Web Store User Data Policy, including the Limited Use requirements.
At the time of installation, Powermeter will present users the option to answer a few multiple-choice questions concerning their gender, country, age, and language. Responding to these questions is facultative and collected answers will only be used to evaluate and compare aggregate click statistics across demographics.
Collected data is securely transmitted (using https protocol) and stored in a database on a secure and protected webserver [..]. Communication between the research team and adherents to the experiment will only occur through in-browser one-way (from the extension to the extension user) pop-up messages. Members of the research team will not be able to individually contact adherent users as they neither know who they are, nor are aware of their contact details.
The processing operations are linked to the organisation, functioning, and management, and follow-up to the Powermeter Research Experiment. Processing operations are limited to what is strictly necessary for the management and functioning of the Powermeter Research Experiment. […] Consequently, by downloading and installing the extension in the Chrome store you are giving us your explicit consent to:
  • Collect, store, and process your numeric installation time-stamp index;
  • Collect, store, and process your clicks associated with different arrangements of content on Google.com, Google.de, Google.ch and Bing.com;
  • Collect, store, and process the answer(s) you will voluntarily disclose concerning your gender, age, location, and language in the onboarding form;
  • Receive pop-up messages for any relevant communication related to the management, functioning, and termination of the Powermeter Research Experiment and to receive the messages hard coded within the extension. Data collection does only begin after Powermeter has been successfully installed.
Your consent can be withdrawn for the future at any time by uninstalling the extension. Your data will be used for the Powermeter Research Experiment only. Your data will not be shared, distributed, or repurposed.
Which personal data do we collect and further process?
No personal data will be collected or processed.
How long do we keep your data?
Your non-personal data will be only kept for the time strictly necessary to fulfil the purpose of collection or further processing.
All data related to the organisation and management of the Powermeter Research Experiment (this includes the information given during the registration, before, and during the experiment) will be deleted after the termination of the research project and within six months at the latest after the last action relating to the research project occurred.
How do we protect and safeguard your data?
All data are collected in electronic format, encrypted as described in Section n. 2, and stored on a secure and protected webserver […]. Access to the server is restricted and controlled.
Who has access to your data and to whom is it disclosed?
Access to your data is provided […] staff responsible for carrying out this processing operation and to other vetted and authorised researchers according to the “need to know” principle. The data we collect will neither be shared, nor transferred to any third party, and will not be repurposed.
What are your rights and how can you exercise them?
As an adherent user you are given the prerogative to interrupt the collection and processing of your data at any time by uninstalling the extension. The withdrawal of your consent will not affect the lawfulness of the processing carried out before you have withdrawn the consent.